A man whose bowing average is 12.4 takes 5 wickets for 26 runs and thereby decreases his average by 0.4 . The number of wickets , taken by him before his last match , is?
1
Expert's answer
2013-02-07T09:30:25-0500
Let the man had taken x wickets with an average of 12.4
Total numbers of runs he gave = 12.4x
Total number of runs he gave after the new spell = 12.4x + 26
New average = (12.4x+26)/(x+5) = 12.4 -0.4 = 12
Solving this equation:
12.4x+26 = 12x+5*12
0.4x = 60 - 26 = 34
x = 34/0.4
x = 85
Answer: Wickets taken by him before his last match = 85